# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ E.2 Identification of Stationary Latent Variables ${\mathbf{z}}_{t}^{s}$
where *** denotes a matrix, and the determinant of this Jacobian matrix is $|\mathbf{J}_{h}|$. Since ${\mathbf{x}}_{t-1}$ do not contain any information of $\hat{{\mathbf{z}}}_{t}$, the right-top element is $\mathbf{0}$. Therefore $p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})\cdot|\mathbf{J}_{h}|$. Dividing both sides of this equation by $p({\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})$ gives
@@
For each $l=n_{e}+1,\cdots n$ and each value of $z_{t-1,l}^{s}$, its partial derivative w.r.t. $z_{t-1,l}^{s}$ is shown as follows
@@
E.3 Identification of Nonstationary Latent Variables ${\mathbf{z}}_{t}^{e}$
Theorem E.4.
(Subspace Identification of the nonstationary latent variables ${\mathbf{z}}^{e}_{t}$.) We follow the data generation process in Figure 2 and Equation (1)-(3), then we make the following assumptions:
  • A3.3.1 ((Smooth and Positive Density:)) The probability density function of latent variables is smooth and positive, i.e. $p({\mathbf{z}}_{t}^{e}|{\mathbf{e}}_{t})>0$ over $\mathcal{Z}_{t}$ and $\mathcal{E}_{t}$.
  • A3.3.2 (Conditional independent:) Conditioned on ${\mathbf{e}}_{t}$, each ${\mathbf{z}}^{e}_{t,i}$ is independent of any other ${\mathbf{z}}_{t,j}^{e}$ for $i,j\in\{1,\cdots,n_{e}\},i\neq j$, i.e., $\log p({\mathbf{z}}_{t}^{e}|{\mathbf{e}}_{t})=\sum_{i=1}^{n_{e}}\log p(z_{t,i}^{e}|{\mathbf{e}}_{t})$.
  • A3.3.3 (Linear independence): For any ${\mathbf{z}}_{t}^{e}\in\mathcal{Z}^{e}_{t}\subseteq\mathbb{R}^{n_{e}}$, there exist $n_{e}+1$ values of ${\mathbf{e}}$, i.e., ${\mathbf{e}}_{j}$ with $j=1,2,..n_{e}+1$, such that these $n_{e}$ vectors $\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})-\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{0})$ are linearly independent, where the vector $\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})$ is defined as follows:
    $\small\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})=\left(\frac{\partial\log p(z_{t,1}^{e}|{\mathbf{e}})}{\partial z_{t,1}^{e}},\cdots,\frac{\partial\log p(z_{t,n_{e}}^{e}|{\mathbf{e}})}{\partial z_{t,n_{e}}^{e}}\right)$ (41)
Then, by learning the data generation process, ${\mathbf{z}}^{e}_{t}$ are subspace identifiable.
@@ E.3 Identification of Nonstationary Latent Variables ${\mathbf{z}}_{t}^{e}$
Assuming that the cross second-order derivative exists (Spantini et al., 2018). Since $p(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})=p(\hat{{\mathbf{z}}}_{t}|\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})p(\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})$ while $p(\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})$ does not involve $\hat{z}_{t,i}$ and $\hat{z}_{t,j}$, the above equality is equivalent to
@@
where *** denotes a matrix, and the determinant of this Jacobian matrix is $|\mathbf{J}_{h}|$. Since ${\mathbf{x}}_{t-1}$ do not contain any information of $\hat{{\mathbf{z}}}_{t}$, the right-top element is $\mathbf{0}$. Therefore $p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})\cdot|\mathbf{J}_{h}|$. Dividing both sides of this equation by $p({\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})$ gives
@@ F.2.1 Dataset Description
  • Exchange (Lai et al., 2018) is the daily exchange rate dataset from of eight foreign countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand, and Singapore ranging from 1990 to.
@@
  • M4 dataset (Makridakis et al., 2020) is a collection of 100,000 time series used for the fourth edition of the Makridakis forecasting Competition with time series of yearly, quarterly, monthly and other (weekly, daily and hourly) data.
@@ F.2.2 Experiment Results on Other Datasets
We further evaluate the proposed method on the traffic and M4 datasets. Experiment results are shown in Table Table 5. According to experiment results of the M4 dataset, which contains the results for yearly, quarterly, and monthly collected univariate marketing data, we can find that the proposed IDEA also outperforms other state-of-the-art deep forecasting models for nonstationary time series forecasting.
@@
models IDEA Koopa SAN DLinear N-Transformer RevIN MICN TimeNets WITRAN
@@ Appendix G Model Efficiency
Following (Liu et al., 2023b) We conduct model efficiency comparasion from three perspectives: forecasting performance, training speed, and memory footprint, which is shown in Figure 6. Compared with other models for nonstationary time-series forecasting, we can find that the proposed IDEA model enjoys the best high model performance and model efficiency, this is because our IDEA is built on MLP-based neural architecture. Compared with other methods like MICN and DLinear, our method achieves a weaker model efficiency, this is because our model needs to model the latent-variable-wise prior.

### v4

@@ E.2 Component-wise Identification of Stationary Latent Variables ${\mathbf{z}}_{t}^{s}$
where * denotes a matrix, and the determinant of this Jacobian matrix is $|\mathbf{J}_{h}|$. Since ${\mathbf{x}}_{t-1}$ do not contain any information of $\hat{{\mathbf{z}}}_{t}$, the right-top element is $\mathbf{0}$. Therefore $p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})\cdot|\mathbf{J}_{h}|$. Dividing both sides of this equation by $p({\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})$ gives
@@
And for $j\in\{n_{e}+1,\cdots,n\}$, the second-order derivative of Equation (57) is
$\small\begin{split}0=\frac{\partial\log p(\hat{z}_{t,i}|\hat{{\mathbf{z}}_{t-1,e_{t}}})}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}&=\sum_{k=1}^{n_{e}}\Big{(}\frac{\partial\log p(z_{t,k}^{e}|e_{t})}{\partial^{2}z_{t,k}^{e}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,j}}+\frac{\partial\log p(z_{t,k}^{e}|e_{t})}{\partial z_{t,k}^{e}}\cdot\frac{\partial^{2}z_{z,k}^{e}}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\Big{)}+\\ \sum_{k=n_{e}+1}^{n}\Big{(}&\frac{\partial\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial^{2}z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}}+\frac{\partial\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\Big{)}+\frac{\partial^{2}\log|\mathbf{J}_{h}|}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\end{split}$ (43)
For each $l=n_{e}+1,\cdots n$ and each value of $z_{t-1,l}$, its partial derivative w.r.t. $z_{t-1,l}$ is shown as follows
$\small\begin{split}0=\frac{\partial^{3}\log p(\hat{z}_{t,i}|\hat{{\mathbf{z}}_{t-1,e_{t}}})}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}\partial z_{t-1,l}}&=\sum_{k=1}^{n_{e}}\Big{(}\frac{\partial^{3}\log p(z_{t,k}^{e}|e_{t})}{\partial^{2}z_{t,k}^{e}\partial z_{t-1,l}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,j}}+\frac{\partial^{2}\log p(z_{t,k}^{e}|e_{t})}{\partial z_{t,k}^{e}\partial z_{t-1,l}}\cdot\frac{\partial^{2}z_{z,k}^{e}}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\Big{)}+\\ \sum_{k=n_{e}+1}^{n}\Big{(}&\frac{\partial^{3}\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial^{2}z_{t,k}^{s}\partial z_{t-1,l}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}}+\frac{\partial^{2}\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}\partial z_{t-1,l}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\Big{)}+\frac{\partial^{3}\log|\mathbf{J}_{h}|}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}\partial z_{t-1,l}}\end{split}$ (44)
Since the distribution $p(z_{t,k}^{e}|e_{t})$ is not influenced by $z_{t-1,l}$, $\frac{\partial^{3}\log p(z_{t,k}^{e}|e_{t})}{\partial^{2}z_{t,k}^{e}\partial z_{t-1,l}}=0$ and $\frac{\partial^{2}\log p(z_{t,k}^{e}|e_{t})}{\partial z_{t,k}^{e}\partial z_{t-1,l}}=0$. Moreover, since $\log|\mathbf{J}_{h}|$ does not depend on $z_{t-1,l}$, $\frac{\partial^{3}\log|\mathbf{J}_{h}|}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}\partial z_{t-1,l}}=0$, and the aforementioned equation can be further rewritten as:
$\small 0=\sum_{k=n_{e}+1}^{n}\Big{(}\frac{\partial^{3}\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial^{2}z_{t,k}^{s}\partial z_{t-1,l}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}}+\frac{\partial^{2}\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}\partial z_{t-1,l}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}\partial\hat{z}_{t,j}}\Big{)}$ (45)
@@
E.3 Component-wise Identification of Nonstationary Latent Variables ${\mathbf{z}}_{t}^{e}$
@@
Assuming that the cross second-order derivative exists (Spantini et al.,, 2018). Since $p(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})=p(\hat{{\mathbf{z}}}_{t}|\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})p(\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})$ while $p(\hat{{\mathbf{z}}}_{t-1},{\mathbf{e}}_{t})$ does not involve $\hat{z}_{t,i}$ and $\hat{z}_{t,j}$, the above equality is equivalent to
@@
where * denotes a matrix, and the determinant of this Jacobian matrix is $|\mathbf{J}_{h}|$. Since ${\mathbf{x}}_{t-1}$ do not contain any information of $\hat{{\mathbf{z}}}_{t}$, the right-top element is $\mathbf{0}$. Therefore $p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})\cdot|\mathbf{J}_{h}|$. Dividing both sides of this equation by $p({\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})$ gives
@@ F.2.1 Dataset Description
  • Exchange (Lai et al.,, 2018) is the daily exchange rate dataset from of eight foreign countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand, and Singapore ranging from 1990 to.
@@
  • M4 dataset (Makridakis et al.,, 2020) is a collection of 100,000 time series used for the fourth edition of the Makridakis forecasting Competition with time series of yearly, quarterly, monthly and other (weekly, daily and hourly) data.
@@ F.2.2 Experiment Results on Other Datasets
We further evaluate the proposed method on the traffic and M4 datasets. Experiment results are shown in Table Table 5. According to experiment results of the M4 dataset, which contains the results for yearly, quarterly, and monthly collected univariate marketing data, we can find that the proposed UDA also outperforms other state-of-the-art deep forecasting models for nonstationary time series forecasting.
@@
models UDA Koopa SAN DLinear N-Transformer RevIN MICN TimeNets WITRAN
@@ F.3 Sensitive Analysis
F.4 Ablation Study
Ablation study of UDA-sh are shown in Table 6. According to the experiment results, we can draw the following conclusions: 1) The performance of the standard UDA and the UDA-Sh are similar, this is because the prediction of $x_{1:t}$ and $x_{t+1:T}$ share the same decoding process. 2) We also find that the performance of UDA is slightly better than that of UDA-Sh in most of the forecasting tasks, reflecting that the model with more parameters may improve the model performance.
Table 6: Experiment results of UDA and UDA-Sh
36-12 72-24 144-48 216-72
Dataset MSE MAE MSE MAE MSE MAE MSE MAE
ECL UDA-Sh 0.123 0.215 0.131 0.257 0.124 0.228 0.132 0.187
UDA 0.114 0.216 0.121 0.22 0.122 0.224 0.131 0.187
ILI UDA-Sh 1.241 0.711 1.69 0.814 1.805 0.866 1.934 0.934
UDA 1.218 0.694 1.68 0.809 1.792 0.869 1.883 0.926
Weather UDA-Sh 0.074 0.093 0.099 0.136 0.121 0.161 0.14 0.193
UDA 0.072 0.09 0.098 0.13 0.115 0.158 0.136 0.187
Exchange UDA-Sh 0.014 0.075 0.024 0.103 0.043 0.143 0.065 0.177
UDA 0.014 0.074 0.023 0.102 0.042 0.141 0.065 0.18
ETTh1 UDA-Sh 0.292 0.344 0.299 0.354 0.355 0.39 0.375 0.395
UDA 0.291 0.345 0.3 0.353 0.338 0.38 0.367 0.388
ETTh2 UDA-Sh 0.14 0.236 0.172 0.26 0.236 0.306 0.288 0.344
UDA 0.141 0.236 0.173 0.26 0.233 0.306 0.262 0.327
@@ Appendix G Model Efficiency
Following (Liu et al., 2023b, ) We conduct model efficiency comparasion from three perspectives: forecasting performance, training speed, and memory footprint, which is shown in Figure 6. Compared with other models for nonstationary time-series forecasting, we can find that the proposed UDA model enjoys the best high model performance and model efficiency, this is because our UDA is built on MLP-based neural architecture. Compared with other methods like MICN and DLinear, our method achieves a weaker model efficiency, this is because our model needs to model the latent-variable-wise prior.